# Beyond Murray’s Law: Non-Universal Branching Exponents from Vessel-Wall Metabolic Costs

### 29 Mar 2026 v2

@@ 3 Mathematical Results
For an asymmetric bifurcation where a parent vessel of flow $Q$ splits into two daughters carrying fractions $fQ$ and (1f)Q(1-f)Q with $f\in(0,1)$, the local branching exponent $\alpha^{*}(Q,f)$ defined by $r^{*}(Q)^{\alpha}=r^{*}(fQ)^{\alpha}+r^{*}((1-f)Q)^{\alpha}$ obeys the same strict boundaries: $(5+p)/2<\alpha^{*}(Q,f)<3$.
@@ 5 Optimal Branching Number
Biological transport networks overwhelmingly favor bifurcations ($N=2$) over higher-order multifurcations ($N>2$). Yet under Murray’s classical cost function, the total energy is completely degenerate with respect to the branching number $N$. The three-term cost function resolves this geometric degeneracy.
@@
For a space-filling fractal network where vessel length scales as $L\propto r$, Murray’s classical cost function ($p=1$) is topologically degenerate, yielding identical total network costs for any branching number $N$. Introducing the empirical sub-linear wall cost ($p<1$) strictly breaks this degeneracy. The competition between the metabolic volume of the vessels (which decreases with $N$) and the steric tissue cost of the junctions (which increases with $N$) strictly forbids star-like topologies ($N\to\infty$) and guarantees the existence of a finite, small optimal branching integer $N^{*}\geq 2$.
@@
Let the network perfuse $M$ terminal units. Under self-similar scaling $r_{g}=r_{0}N^{-2g/(5+p)}$, the metabolic cost of the vessels across generations forms a geometric series. Summing over the network yields the total tube cost: $\mathcal{C}_{tubes}(N)=K_{1}\left[1-N^{-\alpha_{1}}\right]^{-1}$. The series convergence argument fundamentally holds for any finite physiological tree depth $G$: the truncated sum $\sum_{g=0}^{G}N^{-g\alpha_{1}}$ preserves strict monotonicity in $N$ identically to the infinite limit, since each partial sum is a strictly decreasing function of $N$ for $\alpha_{1}>0$. For a finite network depth $G\sim\log_{N}M$, the truncated series yields a macroscopic scaling for the network junction cost:
@@
where $K_{2}>0$. For the empirically observed sub-linear wall scaling $p<1$, the exponent $\alpha_{2}=\frac{1-p}{5+p}$ is strictly positive, making the junction cost strictly increasing with $N$.
@@
Finally, in the classical Murray limit ($p=1$), the exponent $\alpha_{2}=0$, so that $\mathcal{C}_{junctions}(N)$ becomes independent of $N$. Because the tube cost is also $N$-independent under Murray scaling, the total cost becomes completely flat with respect to the branching number, producing a topological degeneracy. Since $H(N)$ is strictly decreasing (established above), the unique critical point of $\mathcal{C}_{\mathrm{tot}}$ on $2,+\infty)$ satisfies $N^{*}_{\mathrm{real}}\in[2,+\infty)$. For all integers $N\geq N^{*}_{\mathrm{real}}$, strict unimodality guarantees $\mathcal{C}_{\mathrm{tot}}(N+1)>\mathcal{C}_{\mathrm{tot}}(N)$, so the integer minimizer is well-defined without further approximation. ∎
@@
Theorem 12 establishes that the metabolic junction penalty strictly bounds the optimal branching number to small finite integers, excluding star-like topologies ($N\to\infty$). However, the slow growth of the junction-cost exponent $(1-p)/(5+p)\approx 0.040$ for $p=0.77$ implies that the purely static cost difference between $N=2$ and $N=3$ is small: junction maintenance alone provides a necessary condition (small $N$) but not a sufficient one (unique $N=2$). A complementary criterion is needed.
@@
Rearranging all $N$-dependent terms to one side defines a function $H(N)$:
@@
Possessing exactly one global minimum and no other local extrema, the function strictly increases for all $N$ beyond its minimum. Thus, $N=2$ is the unique integer minimizer if and only if $\mathcal{C}_{tot}(2)<\mathcal{C}_{tot}(3)$. Evaluating this exact inequality yields the critical threshold:
@@
This structural observation admits a rigorous static completion in the wall-cost-dominated regime. When $\kappa=m_{w}/b\gg 1$, the junction cost overwhelms the tube cost and $\mathcal{C}_{\mathrm{tot}}(N)\approx\mathcal{C}_{\mathrm{junctions}}(N)\propto N^{(1-p)/(5+p)}$, which is strictly increasing in $N$; in this limit, $N=2$ is the unique minimizer of the static cost. For coronary arteries, $\kappa=m_{w}/b\approx 13\gg 1$, so this limit is biologically realized. Together, Theorem 12 (necessary condition: $N<\infty$) and Proposition 14 with the wall-dominated approximation provide a qualitative structural argument for binary branching from wall-cost geometry alone. We note, however, that $\kappa\approx 13$ is pre-asymptotic with respect to the exact threshold $\tau(p)\approx 103$: the wall-dominated limit $\mathcal{C}_{\mathrm{tot}}\approx\mathcal{C}_{\mathrm{junctions}}$ is not yet fully realised at physiological parameters. The asymptotic argument therefore provides a structural insight into the directionality of selection pressure rather than a quantitative demonstration of binary selection—a role filled by Corollary 15. The conclusion is thus robust not through parametric precision but through order-of-magnitude separation: any multiplicative perturbation of $K_{1}$ or $K_{2}$ by a factor less than $10^{2}$ leaves the inequality $K_{2}/K_{1}\ll\tau(p)$ intact.
@@ Results.
Table 2 predicts $r^{*}\in[0.94,1.20]$ mm at $Q=1.3$ mL/s, systematically below the morphometric value $r_{0}=1.5$ mm [6] by 0.30.30.60.6 mm. The present model minimizes the purely passive metabolic cost and therefore predicts the optimal radius absent active tone contributions, not the in-vivo active radius under physiological perfusion pressure.
@@
Since Theorem 4 and Corollary 6 require only that $B,C>0$, the bounds $(5+p)/2<\alpha^{*}(Q)<3$ are preserved exactly under any positive renormalization of $B$. The predicted passive radius therefore represents a testable lower bound on coronary caliber under maximal pharmacological vasodilation, accessible to standard pressure-myograph or hyperemic flow protocols. This discrepancy is physiologically consistent with active vasomotor tone. While our framework predicts the structural (passive) optimum $r^{*}\approx 1.2$ mm, in vivo coronary arteries exhibit significant basal constriction. A vasomotor contraction of approximately 20% would reconcile our theoretical optimum with the higher morphometric values observed in dilated or fixed specimens.
@@ 9 Conclusion
All computation scripts, numerical data, and figure-generation code are openly and unconditionally available at https://github.com/rikymarche-ctrl/vascular-networks-theory under the CC BY 4.0 Licence. No access request is required.

### 27 May 2026 v3

@@ 3 Mathematical Results
For an asymmetric bifurcation where a parent vessel of flow $Q$ splits into two daughters carrying fractions fQfQ and (1f)Q(1-f)Q with $f\in(0,1)$, the local branching exponent $\alpha^{*}(Q,f)$ defined by $r^{*}(Q)^{\alpha}=r^{*}(fQ)^{\alpha}+r^{*}((1-f)Q)^{\alpha}$ obeys the same strict boundaries: $(5+p)/2<\alpha^{*}(Q,f)<3$.
@@ 5 Optimal Branching Number
Biological transport networks overwhelmingly favor bifurcations ($N=2$) over higher-order multifurcations ($N>2$). Yet under Murray’s classical cost function, the total energy is completely degenerate with respect to the branching number NN. The three-term cost function resolves this geometric degeneracy.
@@
For a space-filling fractal network where vessel length scales as $L\propto r$, Murray’s classical cost function ($p=1$) is topologically degenerate, yielding identical total network costs for any branching number NN. Introducing the empirical sub-linear wall cost ($p<1$) strictly breaks this degeneracy. The competition between the metabolic volume of the vessels (which decreases with NN) and the steric tissue cost of the junctions (which increases with NN) strictly forbids star-like topologies ($N\to\infty$) and guarantees the existence of a finite, small optimal branching integer $N^{*}\geq 2$.
@@
Let the network perfuse $M$ terminal units. Under self-similar scaling $r_{g}=r_{0}N^{-2g/(5+p)}$, the metabolic cost of the vessels across generations forms a geometric series. Summing over the network yields the total tube cost: $\mathcal{C}_{tubes}(N)=K_{1}\left[1-N^{-\alpha_{1}}\right]^{-1}$. The series convergence argument fundamentally holds for any finite physiological tree depth $G$: the truncated sum $\sum_{g=0}^{G}N^{-g\alpha_{1}}$ preserves strict monotonicity in NN identically to the infinite limit, since each partial sum is a strictly decreasing function of NN for $\alpha_{1}>0$. For a finite network depth $G\sim\log_{N}M$, the truncated series yields a macroscopic scaling for the network junction cost:
@@
where $K_{2}>0$. For the empirically observed sub-linear wall scaling $p<1$, the exponent $\alpha_{2}=\frac{1-p}{5+p}$ is strictly positive, making the junction cost strictly increasing with NN.
@@
Finally, in the classical Murray limit ($p=1$), the exponent $\alpha_{2}=0$, so that $\mathcal{C}_{junctions}(N)$ becomes independent of NN. Because the tube cost is also NN-independent under Murray scaling, the total cost becomes completely flat with respect to the branching number, producing a topological degeneracy. Since $H(N)$ is strictly decreasing (established above), the unique critical point of $\mathcal{C}_{\mathrm{tot}}$ on $2,+\infty)$ satisfies $N^{*}_{\mathrm{real}}\in[2,+\infty)$. For all integers $N\geq N^{*}_{\mathrm{real}}$, strict unimodality guarantees $\mathcal{C}_{\mathrm{tot}}(N+1)>\mathcal{C}_{\mathrm{tot}}(N)$, so the integer minimizer is well-defined without further approximation. ∎
@@
Theorem 12 establishes that the metabolic junction penalty strictly bounds the optimal branching number to small finite integers, excluding star-like topologies ($N\to\infty$). However, the slow growth of the junction-cost exponent $(1-p)/(5+p)\approx 0.040$ for $p=0.77$ implies that the purely static cost difference between $N=2$ and $N=3$ is small: junction maintenance alone provides a necessary condition (small NN) but not a sufficient one (unique $N=2$). A complementary criterion is needed.
@@
Rearranging all NN-dependent terms to one side defines a function $H(N)$:
@@
Possessing exactly one global minimum and no other local extrema, the function strictly increases for all NN beyond its minimum. Thus, $N=2$ is the unique integer minimizer if and only if $\mathcal{C}_{tot}(2)<\mathcal{C}_{tot}(3)$. Evaluating this exact inequality yields the critical threshold:
@@
This structural observation admits a rigorous static completion in the wall-cost-dominated regime. When $\kappa=m_{w}/b\gg 1$, the junction cost overwhelms the tube cost and $\mathcal{C}_{\mathrm{tot}}(N)\approx\mathcal{C}_{\mathrm{junctions}}(N)\propto N^{(1-p)/(5+p)}$, which is strictly increasing in NN; in this limit, $N=2$ is the unique minimizer of the static cost. For coronary arteries, $\kappa=m_{w}/b\approx 13\gg 1$, so this limit is biologically realized. Together, Theorem 12 (necessary condition: $N<\infty$) and Proposition 14 with the wall-dominated approximation provide a qualitative structural argument for binary branching from wall-cost geometry alone. We note, however, that $\kappa\approx 13$ is pre-asymptotic with respect to the exact threshold $\tau(p)\approx 103$: the wall-dominated limit $\mathcal{C}_{\mathrm{tot}}\approx\mathcal{C}_{\mathrm{junctions}}$ is not yet fully realised at physiological parameters. The asymptotic argument therefore provides a structural insight into the directionality of selection pressure rather than a quantitative demonstration of binary selection—a role filled by Corollary 15. The conclusion is thus robust not through parametric precision but through order-of-magnitude separation: any multiplicative perturbation of $K_{1}$ or $K_{2}$ by a factor less than $10^{2}$ leaves the inequality $K_{2}/K_{1}\ll\tau(p)$ intact.
@@ Results.
Table 2 predicts $r^{*}\in[0.94,1.20]$ $m$ at $Q=1.3$ mL/s, systematically below the morphometric value $r_{0}=1.5$ $m$ [6] by 0.30.30.60.6 mm. The present model minimizes the purely passive metabolic cost and therefore predicts the optimal radius absent active tone contributions, not the in-vivo active radius under physiological perfusion pressure.
@@
Since Theorem 4 and Corollary 6 require only that $B,C>0$, the bounds $(5+p)/2<\alpha^{*}(Q)<3$ are preserved exactly under any positive renormalization of $B$. The predicted passive radius therefore represents a testable lower bound on coronary caliber under maximal pharmacological vasodilation, accessible to standard pressure-myograph or hyperemic flow protocols. This discrepancy is physiologically consistent with active vasomotor tone. While our framework predicts the structural (passive) optimum $r^{*}\approx 1.2$ $m$, in vivo coronary arteries exhibit significant basal constriction. A vasomotor contraction of approximately 20% would reconcile our theoretical optimum with the higher morphometric values observed in dilated or fixed specimens.
@@ 9 Conclusion
All computation scripts, numerical data, and figure-generation code are openly and unconditionally available at https://github.com/rikymarche-ctrl/vascular-networks-theory under the $C$ BY 4.0 Licence. No access request is required.